OpenStudy (anonymous):

solve please: x^4+4x^2-12=0

OpenStudy (acxbox22):

what are 2 factors that multiply to -12 and add up to 4

OpenStudy (acxbox22):

hint:put those factors here (x^2- )(x^2+ )

OpenStudy (acxbox22):

set equal to zero and solve

OpenStudy (anonymous):

this is factorable. put it into this format: (x+?)(x+?). the numbers should be two numbers that multiply to -12 (a*c) and those same two numbers that add up to 4 (b) if ax^2+bx+c=0.

OpenStudy (anonymous):

i'll give you a clue: one of the numbers is positive 6.

OpenStudy (anonymous):

after you put it into (x+k)(x+l) (where k and l are the 2 numbers), do x+k=0 and x+l=0 and there are your answers.